OpenStudy (anonymous):

57 – 12 ÷ 4 • 3 help pls

OpenStudy (anonymous):

PEMDAS: Parentheses, Exponents, Multiplication, Division, Addition, Subtraction. 57 - 12 / 4(3) 57 - 12 / 12 57 - 1 56

OpenStudy (rina.r):

its BIMDAS not pemdas

OpenStudy (anonymous):

so whats the answer?

OpenStudy (rina.r):

is it bract or point?

OpenStudy (rina.r):

*bracket

OpenStudy (anonymous):

is says Simplify the expression thats all it says

OpenStudy (rina.r):

but a point and a bracket can change teh answer so all i am saying is that is it a bracket after 4 or point?

OpenStudy (anonymous):

it's a symbol meaning multiplication.

OpenStudy (rina.r):

ohh ok

OpenStudy (anonymous):

o.O

OpenStudy (rina.r):

57 – 12 ÷ 4 x 3 57-12/ 12 1 57-1 so 56 is the right answer.

OpenStudy (anonymous):

Ok thank you

OpenStudy (rina.r):

12 divide by 12 gives u 1 an dthen u subtract from 57 gives u 56 by using BIMDAS rule.

OpenStudy (rina.r):

welcome

OpenStudy (anonymous):

i have another question 7 × [8 + (16 – 9)] ÷ 5

OpenStudy (rina.r):

7x(8+(16-9)/5 7x(8+7)/5 7x15/5 105/5 =21

OpenStudy (anonymous):

Oh! i see

OpenStudy (rina.r):

yeah u see got it?

OpenStudy (anonymous):

use the bodmas rule that is(bracket off division multiplication addition substraction)

OpenStudy (anonymous):

Yes I see lol

OpenStudy (rina.r):

guys its BIMDSA not bodmas

OpenStudy (anonymous):

i have 3 more questions

OpenStudy (rina.r):

okkk go head

OpenStudy (anonymous):

Which property justifies this statement? If x + y = z and z = x + 4, then x + y = x + 4. A. symmetric B. associative C. distributive D. transitive

OpenStudy (anonymous):

oh sorry lol i dont know how to do that

OpenStudy (anonymous):

if i knew i would give you 10000000 medals

OpenStudy (anonymous):

Which property justifies this statement? If x + y = z and z = x + 4, then x + y = x + 4. A. symmetric B. associative C. distributive D. transitive

OpenStudy (anonymous):

lool :D ......i will show u ....just press on the medal sign :)

OpenStudy (anonymous):

do you mean press best respond

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

oh i see

OpenStudy (rina.r):

its B. associative

OpenStudy (anonymous):

thanks

OpenStudy (rina.r):

welcome

OpenStudy (anonymous):

Translate the word phrase into a math expression. six less than eight times a number A. 6n – 8 B. 6 – 8n C. 8n – 6 D. 6 < 8n